OpenStudy (anonymous):

Integration of 2/1-t^2 Steps please.

OpenStudy (anonymous):

=1/(1-t) + 1/(1+t) = ln(1+t) -ln(1-t) +c =ln((1+t)/(1-t)) +c

OpenStudy (anonymous):

Its a direct formula actually...